OpenStudy (anonymous):

represent this polar equation: r=a(1+cosine of theta) in a cartesian form

OpenStudy (anonymous):

use x^2 + y^2 = r^2 ans tan theta = sin theta / cos theta = y / x

OpenStudy (anonymous):

\(\large r=a(1+cos\theta) \) square both sides to get \(\large r^2=a^2(1+cos\theta)^2 \) then expand that right side then you can make substitutions using these relations: \(\large x=rcos\theta \) and \(\large y=rsin\theta \) and also \(\large x^2+y^2=r^2 \)

OpenStudy (anonymous):

@dpalnc...i got stuck. the degree of the equation i'm getting is scary. very scary. can u help me out with the final ans so i could compare

OpenStudy (anonymous):

@Aparna2...yea i started with dat already but then it got very ugly along the line with an equation with very high degrees

OpenStudy (anonymous):

It is just some algebra, nothing higher than quadratic.....